# Poisoning with A Pill: Circumventing Detection in Federated Learning

### v1

@@ 6.1 Evaluation Settings
Model, Dataset, and Hyper-Parameters. In our experiments, we employ a four-layer Convolutional Neural Network (CNN) and a simplified version of AlexNet [54]. The structures of the models and their corresponding pill blueprints are detailed in Appendix B. We evaluate our method on three widely-used datasets: MNIST [55], Fashion-MNIST [56], and CIFAR-10 [57]. We use the CNN model on MNIST and Fashion-MNIST datasets, and the AlexNet [54] on CIFAR-10 dataset. Each experiment is repeated five times to ensure reliability, with the mean and standard deviation (std) of the results reported.
TABLE III: Error rates under cross-silo setting using “approximate max pill search” (20% malicious clients) on Fashion-MNIST dataset.
Data Distribution IID Non-IID
$\pm$0.003 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.002 $\pm$0.002 $\pm$0.002 $\pm$0.003 $\pm$0.004 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.003
Sign-flipping Attack 0.943 0.114 0.108 0.126 0.136 0.116 0.118 0.917 0.126 0.117 0.132 0.152 0.124 0.127
$\pm$0.023 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.002 $\pm$0.001 $\pm$0.003 $\pm$0.020 $\pm$0.004 $\pm$0.002 $\pm$0.003 $\pm$0.006 $\pm$0.003 $\pm$0.003
0.667 0.115 0.764 0.379 0.523 0.314 0.646 0.543 0.122 0.754 0.430 0.522 0.311 0.688
+ Poison Pill $\pm$0.089 $\pm$0.004 $\pm$0.049 $\pm$0.104 $\pm$0.091 $\pm$0.018 $\pm$0.061 $\pm$0.150 $\pm$0.006 $\pm$0.129 $\pm$0.057 $\pm$0.038 $\pm$0.038 $\pm$0.067
$\pm$0.010 $\pm$0.003 $\pm$0.002 $\pm$0.006 $\pm$0.006 $\pm$0.003 $\pm$0.001 $\pm$0.022 $\pm$0.005 $\pm$0.018 $\pm$0.004 $\pm$0.008 $\pm$0.009 $\pm$0.003
0.618 0.576 0.638 0.284 0.453 0.219 0.115 0.668 0.517 0.687 0.292 0.473 0.223 0.222
+ Poison Pill $\pm$0.071 $\pm$0.057 $\pm$0.041 $\pm$0.040 $\pm$0.091 $\pm$0.010 $\pm$0.003 $\pm$0.033 $\pm$0.038 $\pm$0.036 $\pm$0.047 $\pm$0.047 $\pm$0.016 $\pm$0.128
$\pm$0.002 $\pm$0.003 $\pm$0.022 $\pm$0.009 $\pm$0.008 $\pm$0.003 $\pm$0.087 $\pm$0.004 $\pm$0.003 $\pm$0.059 $\pm$0.003 $\pm$0.005 $\pm$0.003 $\pm$0.030
0.735 0.155 0.715 0.422 0.578 0.310 0.637 0.716 0.151 0.737 0.468 0.730 0.334 0.690
+ Poison Pill $\pm$0.032 $\pm$0.032 $\pm$0.132 $\pm$0.046 $\pm$0.057 $\pm$0.009 $\pm$0.074 $\pm$0.104 $\pm$0.004 $\pm$0.078 $\pm$0.017 $\pm$0.168 $\pm$0.031 $\pm$0.079
$\pm$0.008 $\pm$0.002 $\pm$0.029 $\pm$0.015 $\pm$0.009 $\pm$0.009 $\pm$0.052 $\pm$0.026 $\pm$0.015 $\pm$0.050 $\pm$0.080 $\pm$0.012 $\pm$0.007 $\pm$0.049
0.702 0.303 0.668 0.327 0.514 0.314 0.778 0.629 0.320 0.612 0.406 0.547 0.376 0.822
+ Poison Pill $\pm$0.114 $\pm$0.201 $\pm$0.116 $\pm$0.074 $\pm$0.053 $\pm$0.047 $\pm$0.063 $\pm$0.114 $\pm$0.115 $\pm$0.040 $\pm$0.065 $\pm$0.072 $\pm$0.119 $\pm$0.036
@@
IID and Non-IID Data Settings. Our method was assessed under both IID and non-IID data distributions to understand its performance across data heterogeneity. For IID data setting, we uniformly split all the training data into $K$ shards, and distribute each shard to a random client. For non-IID data setting, we utilize the non-IID degree $p$ as defined in prior studies [4, 30]. A higher $p$ indicates greater data heterogeneity among the clients. Specifically, when $p=0.1$, the data configuration is essentially IID. We set $p=0.5$ to to intensify the non-IID condition, under which we create and allocate $K$ non-IID data shards to all the clients, simulating a more realistic and challenging FL environment. Given that FLTrust necessitates a root dataset at the server, we select this dataset first from the available training data. Subsequently, we distribute the remaining data among the clients according to the aforementioned IID and non-IID rules. This approach ensures that there is no overlap between server’s data and client’s data.
@@
Configurations of Dynamic Patterns in Our Method. As outlined in Section 5.2, we design six dynamic patterns for the pill search. We systematically evaluate all six patterns and present the results of the most effective strategy.
@@
Evaluation Metrics. We use error rates – defined as the proportion of incorrect predictions – to evaluate attack effectiveness. Given that the model poisoning attacks discussed are all untargeted, higher error rates indicate more effective attacks. To assess the stealthiness of our method in delivering malicious updates, we employ two metrics: 1) cosine similarity score: measures alignment with the server’s model update in FLTrust; 2) distance score: used in Multi-Krum to evaluate the closeness of poisoned updates to benign updates.
TABLE IV: Error rates under cross-silo setting using “approximate max pill search” (10% malicious clients) on Fashion-MNIST dataset.
Data Distribution IID Non-IID
$\pm$ 0.003 $\pm$0.003 $\pm$0.003 $\pm$0.004 $\pm$0.001 $\pm$0.002 $\pm$0.002 $\pm$0.002 $\pm$0.003 $\pm$0.001 $\pm$0.002 $\pm$0.005 $\pm$0.002 $\pm$0.004
Sign-flipping Attack 0.964 0.109 0.108 0.110 0.130 0.108 0.117 0.909 0.119 0.114 0.119 0.144 0.120 0.125
$\pm$0.017 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.005 $\pm$0.001 $\pm$0.005 $\pm$0.045 $\pm$0.002 $\pm$0.003 $\pm$0.002 $\pm$0.004 $\pm$0.004 $\pm$0.002
0.320 0.116 0.162 0.151 0.323 0.148 0.699 0.269 0.120 0.239 0.164 0.364 0.168 0.242
+ Poison Pill $\pm$0.080 $\pm$0.007 $\pm$0.027 $\pm$0.010 $\pm$0.029 $\pm$0.007 $\pm$0.082 $\pm$0.174 $\pm$0.003 $\pm$0.101 $\pm$0.013 $\pm$0.031 $\pm$0.012 $\pm$0.198
$\pm$0.002 $\pm$0.005 $\pm$0.004 $\pm$0.003 $\pm$0.004 $\pm$0.003 $\pm$0.001 $\pm$0.003 $\pm$0.006 $\pm$0.001 $\pm$0.004 $\pm$0.005 $\pm$0.002 $\pm$0.002
0.508 0.139 0.334 0.126 0.284 0.127 0.120 0.528 0.148 0.455 0.143 0.287 0.146 0.136
+ Poison Pill $\pm$0.128 $\pm$0.012 $\pm$0.120 $\pm$0.006 $\pm$0.040 $\pm$0.004 $\pm$0.003 $\pm$0.051 $\pm$0.018 $\pm$0.151 $\pm$0.003 $\pm$0.023 $\pm$0.004 $\pm$0.012
$\pm$0.004 $\pm$0.005 $\pm$0.001 $\pm$0.002 $\pm$0.004 $\pm$0.004 $\pm$0.134 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.001 $\pm$0.005 $\pm$0.003 $\pm$0.352
0.183 0.118 0.283 0.161 0.362 0.146 0.631 0.428 0.127 0.280 0.187 0.415 0.182 0.704
+ Poison Pill $\pm$0.039 $\pm$0.003 $\pm$0.210 $\pm$0.015 $\pm$0.027 $\pm$0.008 $\pm$0.089 $\pm$0.190 $\pm$0.005 $\pm$0.064 $\pm$0.012 $\pm$0.057 $\pm$0.009 $\pm$0.091
$\pm$0.004 $\pm$0.004 $\pm$0.002 $\pm$0.005 $\pm$0.009 $\pm$0.008 $\pm$0.004 $\pm$0.003 $\pm$0.012 $\pm$0.003 $\pm$0.007 $\pm$0.004 $\pm$0.004 $\pm$0.002
0.439 0.129 0.361 0.136 0.343 0.150 0.715 0.521 0.136 0.339 0.153 0.368 0.184 0.335
+ Poison Pill $\pm$0.140 $\pm$0.009 $\pm$0.245 $\pm$0.015 $\pm$0.032 $\pm$0.006 $\pm$0.096 $\pm$0.073 $\pm$0.011 $\pm$0.202 $\pm$0.009 $\pm$0.048 $\pm$0.017 $\pm$0.185
TABLE V: Error rates under cross-silo setting using “approximate max pill search” (20% malicious clients) on CIFAR-10 dataset.
Data Distribution IID Non-IID
Attack FedAvg FLTrust MKrum Bulyan Median Trim DnC FLD Flame FedAvg FLTrust MKrum Bulyan Median Trim DnC FLD Flame
No Attack 0.488 0.480 0.507 0.469 0.551 0.456 0.445 0.494 0.491 0.486 0.474 0.499 0.498 0.581 0.502 0.463 0.506 0.532
Sign-flipping Attack 0.898 0.479 0.580 0.539 0.621 0.461 0.468 0.497 0.509 0.905 0.514 0.511 0.622 0.658 0.573 0.502 0.603 0.533
+ Poison Pill 0.739 0.880 0.929 0.694 0.707 0.699 0.536 0.899 0.706 0.879 0.861 0.898 0.677 0.766 0.688 0.566 0.900 0.675
Trim Attack 0.482 0.509 0.489 0.536 0.623 0.514 0.456 0.459 0.501 0.571 0.493 0.608 0.595 0.653 0.549 0.481 0.482 0.506
+ Poison Pill 0.853 0.877 0.883 0.654 0.674 0.662 0.513 0.899 0.542 0.890 0.862 0.906 0.772 0.688 0.639 0.518 0.893 0.621
Krum Attack 0.473 0.541 0.471 0.568 0.540 0.510 0.455 0.802 0.500 0.485 0.506 0.497 0.522 0.647 0.519 0.481 0.899 0.501
+ Poison Pill 0.701 0.896 0.900 0.765 0.756 0.643 0.529 0.890 0.872 0.724 0.849 0.900 0.675 0.748 0.647 0.580 0.885 0.873
Min-Max Attack 0.450 0.504 0.469 0.507 0.579 0.465 0.514 0.525 0.525 0.478 0.502 0.493 0.568 0.636 0.603 0.478 0.482 0.488
+ Poison Pill 0.752 0.712 0.902 0.775 0.802 0.640 0.545 0.902 0.811 0.661 0.646 0.886 0.674 0.783 0.661 0.527 0.907 0.799
Figure 8: Comparison of error rates between original poisoning attacks and our method with two different pill search methods.
6.2 Augmentation Effectiveness
In this section, we present a comprehensive analysis of our method’s augmentation effectiveness on Fashion-MNIST dataset within a 50-client cross-silo FL system, in which 20% clients are malicious. We evaluate our method on both IID data and non-IID data. Our method successfully augments all the baseline attacks with more than $0.25$ average error rate increase, showing our method’s effectiveness and high compatibility.
Results on IID Data. The error rates of four baseline FL poisoning attacks, with and without our method, are shown in the left half of Table III. Our method enhances the error rates of the existing poisoning attacks in $23$ out of $28$ scenarios, against FedAvg and five defenses. The maximum increase in error rate is $0.658$, and the average increase reaches $0.274$. This substantial elevation from the attack-free baseline error rate of $0.109$ underscores our method’s capability to significantly compromise existing defenses’ integrity. Following are individual improvements of our method on different baseline attacks:
@@ 6.2 Augmentation Effectiveness
Results on Non-IID Data. Evaluations on non-IID data further validate the effectiveness of our method, demonstrating its superiority in $23$ out of $28$ cases. The highest error rate increase reaches $0.637$, with an average increase of $0.281$. Although there is a slight reduction in maximal error rate increase in the non-IID data setting, these results still demonstrate our method’s ability to effectively enhance attacks in more complex and heterogeneous data environments. The detailed improvements for a specific attack are shown as follows:
@@
All attacks augmented by our method can bypass all baseline defenses, including FLTrust and FLDetector, with the exception of the sign-flipping attack. Notably, the Min-Max attack demonstrates superior effectiveness in non-IID data settings, achieving significant improvements compared to its performance on IID data. Other attacks also exhibit similar error rate improvements relative to their results on IID data, indicating that our method maintains its robustness and effectiveness in more complex data environments.
6.3 Stealthiness Analysis
To further analyze the performance of our method, we analyze its stealthiness during the training process of the FL system, focusing on how our method influences the distance scores and cosine similarity scores of exisitng FL poisoning attacks. The results indicate that our method can make malicious clients appear as benign or even more “benign” than genuine benign clients. This significant increase in stealthiness is a result of the pill design with distance-based adjustment and similarity-based adjustment techniques in our method.
Distance Score Analysis. Figure 5 compares the average distance scores of benign and malicious clients (with and without our method) across four baseline model poisoning attacks. The distance scores when using our method closely match or are even identical to those of benign clients throughout the entire training process. In contrast, original attacks like the Trim and sign-flipping attacks display distance scores that were significantly higher or lower than those of benign updates, indicating either detected by Multi-Krum (higher scores) or underutilized attack capacities (lower scores). Our method also has lower distance score variance in the early FL training period, representing that our method provides more steady attack efficacy in the FL’s critical training period [58, 59] by fully utilizing the attack capacities while being undetected. More details are shown in Appendix D.
@@ 6.3 Stealthiness Analysis
6.4 Generality Analysis
In this section, we further discuss the generality of our method from four perspectives: malicious client proportion, client participation frequency, datasets & model architectures, and pill search algorithm. The results indicate that our method maintains its augmentation effectiveness consistently, even as these conditions change, demonstrating its reliability and wide applicability in augmenting FL poisoning attacks.
Impact of The Malicious Client’s Proportion. We first assess the effectiveness of our method in both IID and non-IID cross-silo FL system with only $10$% of clients compromised, as shown in Table IV. This setup revealed that all baseline model poisoning attacks yield lower error rates on the global model compared with scenarios with $20$% compromised clients. While the increase in error rates was less than those in the $20$% compromised client scenario, our method still effectively raises the global model’s error rates in $25$/$26$ out of $28$ cases (IID/non-IID setting). The maximum increase in error rates reached $0.403$, with an average increase of $0.144$. This average is notably higher (>$2$x higher) than the error rates observed in attack-free FL conditions. Specifically, our method help sign-flipping/Trim/Krum/Min-Max attacks achieve an average error rate increase of $0.133$/$0.094$/$0.136$/$0.209$. More detailed results are presented in Appendix G.
Impact of The Client Participation Frequency. We then extend the evaluation of our method to a cross-device FL system, where only $40$% of clients are selected for participation in each communication round. This setup results in less frequent participation from each client and a fluctuating proportion of malicious clients across different rounds. The maximum error rate increase with our method is $0.639$, with an average increase across different attacks and defenses of $0.279$. These results are consistent with those from the cross-silo FL system, underscoring our method’s effectiveness and generality across different FL configurations. This evaluation demonstrates our method’s robust performance and adaptability, not only in a controlled cross-silo environment but also under the more various conditions in cross-device FL systems. More details are presented in Appendix F.
Impact of The Datasets and Model Architectures. Following the evaluation with the Fashion-MNIST dataset, we test our method on the MNIST and CIFAR-10 dataset, employing the four-layer CNN model and the AlexNet model to further verify our method’s generality across different datasets. The collective results show that our method performs even better with larger datasets or more complex machine learning models. This trend confirms the generality of our method by revealing its capability to maintain consistent performance enhancements regardless of the dataset or model complexity involved. Specifically, our method help all the four baseline attacks bypass all the nine baselines on CIFAR-10 dataset, achieving $0.288$ error rate increase on average, presented in Table V. More detailed results on MNIST dataset are shown in Appendix E.
@@ 6.4 Generality Analysis
Impact of The Pill Search Algorithm. We conduct a final evaluation to assess the importance and effectiveness of the "approximate max pill search" algorithm used in our method. This is contrasted against a newly devised "approximate min pill search" algorithm, which targets the least important parameters within the target model. Figure 8 illustrates the error rates achieved by the "approximate max pill search", the "approximate min pill search", and the original model poisoning attacks. The "approximate max pill search" algorithm outperforms the "approximate min pill search" in $41$ out of $56$ cases (approximately $73$%), underscoring its effectiveness in leveraging the most influential parameters to enhance attack impacts. Despite its lower efficacy, the "approximate min pill search" still manage to surpass the original attacks in $41$ out of $56$ cases (approximately $73$%). This demonstrates the generality of our method across different pill search algorithms.
7 Discussion
7.1 Our Method against Possible Adaptive Defense
@@
To further evaluate the robustness of our method when defenses are aware of the attack strategies (white-box scenario), we develop an adaptive defense named DSTrust, which enhances the FLTrust’s mechanism. DSTrust incorporates both distance and cosine similarity scores into a unified trust score calculation, directly countering our method’s two-step adjustment approach. The round-$t$ trust score of client $i$ in DSTrust is calculated as follows:
$TS_{i}=ReLU(\frac{\mathop{cos}(\Delta\boldsymbol{g}_{t}^{(i)},\Delta\boldsymbol{g}_{t}^{s})}{||\Delta\boldsymbol{g}_{t}^{(i)}-\Delta\boldsymbol{g}_{t}^{s}||}),$ (3)
@@
TABLE VI: Error rates under cross-silo setting against the new adaptive defense – DSTrust – with and without our method on Fashion-MNIST dataset (20% malicious clients).
Data Distribution IID Non-IID
@@
 Poison Pill
No Attack 0.108 0.116
@@
 Attack 0.111 0.129 0.110 0.131
Trim Attack 0.109 0.629 0.115 0.630
Krum Attack 0.111 0.140 0.120 0.128
Min-Max Attack 0.127 0.167 0.143 0.327
Table VI details the error rates for four baseline FL poisoning attacks both with and without our method against the DSTrust defense on the Fashion-MNIST dataset within a 50-client FL system, where 20% of clients are malicious. These tests were conducted under both IID and non-IID data environments. DSTrust effectively neutralizes the four baseline poisoning attacks when our method is not applied, highlighting its robustness as a defense mechanism. Despite DSTrust’s integration of both cosine similarity and distance metrics in its defense strategy, it fails to counteract the augmented attacks when our method is employed. Notably, our method achieves a maximum error rate increase of $0.521$, and an average error rate increase of $0.173$ across all $8$ test scenarios. These results demonstrate that merely understanding the adjustment strategies of our method, and subsequently integrating corresponding defense metrics, does not fundamentally negate the effectiveness of our method. Despite the adaptive defense’s attempt to incorporate both cosine similarity and distance metrics into DSTrust, it remains insufficient to thwart the enhanced capabilities of our method.
7.2 Limitations and Future Work

### v2

@@ Appendix H Additional Experimental Configurations
Model, Dataset, and Hyper-Parameters. In our experiments, we employ a four-layer Convolutional Neural Network (CNN) and a simplified version of AlexNet (alexnet). The structures of the models and their corresponding pill blueprints are detailed in Appendix D. We evaluate our method on three widely-used datasets: MNIST (lecun1998mnist), Fashion-MNIST (fashionmnist), and CIFAR-10 (cifar). We use the CNN model on MNIST and Fashion-MNIST datasets, and the AlexNet (alexnet) on CIFAR-10 dataset. Each experiment is repeated five times to ensure reliability, with the mean and standard deviation (std) of the results reported.
@@
IID and Non-IID Data Settings. Our method was assessed under both IID and non-IID data distributions to understand its performance across data heterogeneity. For IID data setting, we uniformly split all the training data into $K$ shards, and distribute each shard to a random client. For non-IID data setting, we utilize the non-IID degree $p$ as defined in prior studies (fang2020local; guo2021siren). A higher $p$ indicates greater data heterogeneity among the clients. Specifically, when $p=0.1$, the data configuration is essentially IID. We set $p=0.5$ to to intensify the non-IID condition, under which we create and allocate $K$ non-IID data shards to all the clients, simulating a more realistic and challenging FL environment. Given that FLTrust necessitates a root dataset at the server, we select this dataset first from the available training data. Subsequently, we distribute the remaining data among the clients according to the aforementioned IID and non-IID rules. This approach ensures that there is no overlap between server’s data and client’s data.
@@
Configurations of Dynamic Patterns. As outlined in Section 4.2, we design six dynamic patterns for the pill search. We systematically evaluate all six patterns and present the results of the most effective strategy.
@@
Evaluation Metrics. We use error rates – defined as the proportion of incorrect predictions – to evaluate attack effectiveness. Given that the model poisoning attacks discussed are all untargeted, higher error rates indicate more effective attacks. To assess the stealthiness of our method in delivering malicious updates, we employ two metrics: 1) cosine similarity score, measuring alignment with the server’s model update in FLTrust; 2) distance score, used in Multi-Krum to evaluate the closeness of poisoned updates to benign updates.
Appendix I Additional Augmentation Performance Analysis
Following are individual improvements of our method on different baseline attacks in IID data setting:
@@ Appendix I Additional Augmentation Performance Analysis
Similarly, the detailed improvements for a specific attack in the non-IID data setting shown as follows:
@@
Data Distribution IID Non-IID
Sign-Flipping 0.934 0.059 0.038 0.055 0.055 0.036 0.025 0.886 0.073 0.041 0.052 0.059 0.041 0.026
[…]
Table 8: Error rates on MNIST under cross-silo setting with 20% malicious clients in the 50-client FL system.
Data Distribution IID Non-IID
Sign-Flipping 0.940 0.116 0.110 0.128 0.165 0.121 0.905 0.124 0.118 0.136 0.184 0.134
[…]
Table 9: Error rates on Fashion-MNIST under cross-device setting with 20% malicious clients in 50-client FL system.
Data Distribution IID Non-IID
Sign-Flipping 0.929 0.111 0.108 0.111 0.153 0.117 0.902 0.118 0.115 0.120 0.175 0.134
[…]
Table 10: Error rates on Fashion-MNIST under cross-device setting with 10% malicious clients in 50-client FL system.
Appendix J Additional Stealthiness Analysis
Distance Score Analysis. Figure 4 compares the average distance scores of benign and malicious clients (with and without our method) across four baseline model poisoning attacks. The distance scores when using our method closely match or are even identical to those of benign clients throughout the entire training process. In contrast, original attacks like the Trim and sign-flipping attacks display distance scores that were significantly higher or lower than those of benign updates, indicating either detected by Multi-Krum (higher scores) or underutilized attack capacities (lower scores). Our method also has a lower distance score variance in the early FL training period, representing that our method provides more steady attack efficacy in the FL’s critical training period (yan2023defl; yan2022seizing) by fully utilizing the attack capacities while being undetected. Additionally, our method also achieves two more improvements. First, our method causes the global model to degrade earlier compared to the original attacks, further demonstrating the effectiveness of our augmentation. Second, our method significantly increases the discrepancy between benign client updates as the communication rounds increase. While original attacks can bypass detection in some cases, the discrepancy between benign client updates remains steady, illustrating the lower impact of malicious clients. In contrast, our method consistently increases the discrepancy among benign clients, highlighting its penetrating effectiveness in its influence on benign clients’ local training.
@@
Appendix K Additional Details on MNIST and CIFAR-10 Dataset
The detailed results on MNIST and CIFAR-10 datasets are presented in Table 8 (MNIST dataset) and Table 3 (CIFAR-10 dataset), respectively.
On CIFAR-10 dataset, our method helps existing FL poisoning attacks outperform their original versions in $71$ of the $72$ scenarios, with an average error rate increase of over $0.288$. Specifically, our method facilitates at least a $0.212$ increase in error rates against FLTrust, outperforming the results in the same settings on the Fashion-MNIST dataset.
Appendix L Additional Results in Cross-device FL System
After evaluating our method in the 50-client cross-silo FL system, we further test it in the 50-client cross-device FL system. Table 9 presents the error rates under the cross-device FL setting using the “approximate max pill search” algorithm on both IID and non-IID data. We report the highest error rates among the results of six dynamic patterns, with the malicious client proportion set to $20\%$. Since FLD is not typically designed for cross-device systems, we do not test it in this setting.
Results on IID Data. The highest error rate improvement with our method achieves $0.639$, and the average error rate increase with our method reaches $0.279$. With our method, existing model poisoning attacks outperform their original versions in $22$ out of the $24$ cases. The highest error rate improvement for the sign-flipping attack is $0.639$, with an average error rate increase of $0.279$. For the Trim attack and Krum attack, the highest error rate increases are $0.469$ and $0.568$, with average error rate increases of $0.264$ and $0.302$, respectively. For the Min-Max attack, the highest error rate increase reaches $0.505$, with an average increase of $0.272$. These improvements are consistent with the error rates observed under the cross-silo FL setting using the ”approximate max pill search” algorithm on IID data.
Results on non-IID Data. As for the results on non-IID data, the highest error rate improvement with our method achieves $0.546$, and the average error rate increase with our method reaches $0.273$. By using our method, existing model poisoning attacks outperform their original versions in 21 out of the 24 cases. The highest error rate improvement for the sign-flipping attack is $0.547$, with an average error rate increase of $0.282$. For the Trim attack and Krum attack, the highest error rate rises are $0.451$ and $0.546$, with an average error rate rise of $0.312$ and $0.278$, respectively. For the Min-Max attack, the highest error rate increase reaches $0.479$, with an average increase of $0.201$. These improvements are also aligned with the error rates observed under the cross-silo FL setting using the max subnetwork searching algorithm on non-IID data.
Appendix M Additional Results with Fewer Malicious Clients
We also test the error rate improvement of our method in both the IID and non-IID cross-device FL systems, with only $10\%$ malicious clients. The experimental results are shown in Table 10.
Appendix N Impact of The Pill Search Algorithm
Figure 8: Comparison of error rates between original poisoning attacks and attacks enhanced by our method using two different pill search methods.
@@
We conduct a final evaluation to assess the importance and effectiveness of the ”approximate max pill search” algorithm used in our method. This is contrasted against a newly devised ”approximate min pill search” algorithm, which targets the least important parameters within the target model. Figure 8 illustrates the error rates achieved by the ”approximate max pill search”, the ”approximate min pill search”, and the original model poisoning attacks. The ”approximate max pill search” algorithm outperforms the ”approximate min pill search” in $41$ out of $56$ cases (approximately 7373%), underscoring its effectiveness in leveraging the most influential parameters to enhance attack impacts. Despite its lower efficacy, the ”approximate min pill search” still manages to surpass the original attacks in $41$ out of $56$ cases (approximately 7373%). This demonstrates the generality of our method across different pill search algorithms.
Appendix O Our Method Against Adaptive Defense
@@
We develop an adaptive defense named DSTrust, which enhances the FLTrust’s mechanism. DSTrust incorporates both distance and cosine similarity scores into a unified trust score calculation, directly countering our method’s two-step adjustment approach. The round-$t$ trust score of client $i$ in DSTrust is calculated as follows:
[…]
@@
Distribution IID Non-IID
Attack       Poison Pillw/o      Poison Pillw/       Poison Pillw/o      Poison Pillw/
No Attack 0.108 0.116
@@
Table 11: Error rates under cross-silo setting against the new adaptive defense – DSTrust – with and without our method on Fashion-MNIST dataset (20% malicious clients).
Data Distribution IID Non-IID
Sign-Flipping 0.943 0.114 0.108 0.126 0.136 0.116 0.118 0.917 0.126 0.117 0.132 0.152 0.124 0.127
          w/ DistAdjust       w/o SimAdjust+ Poison Pill        0.311 0.096 0.219 0.413 0.354 0.381 0.726 0.372 0.111 0.224 0.373 0.298 0.448 0.722
          w/o DistAdjust      w/ SimAdjust+ Poison Pill       0.826 0.110 0.118 0.142 0.898 0.464 0.695 0.849 0.117 0.133 0.161 0.870 0.491 0.634
@@
          w/ DistAdjust       w/o SimAdjust+ Poison Pill        0.317 0.105 0.364 0.247 0.368 0.136 0.208 0.219 0.111 0.600 0.456 0.295 0.212 0.105
          w/o DistAdjust      w/ SimAdjust+ Poison Pill       0.554 0.104 0.122 0.108 0.429 0.284 0.119 0.744 0.110 0.134 0.126 0.898 0.478 0.109
[…]
Table 12: Ablation study on the necessity of both SimAdjust and DistAdjust in our method on Fashion-MNIST dataset with 20% malicious clients in the 50-client FL system.
Data Distribution IID Non-IID
Label-Flipping 0.937 0.110 0.103 0.115 0.125 0.122 0.097 0.767 0.118 0.108 0.128 0.126 0.125 0.104
[…]
Table 13: Error rates of Label-flipping Attack on Fashion-MNIST under cross-silo setting with 20% malicious clients in the 50-client FL system.
Table 11 details the error rates for four baseline FL poisoning attacks both with and without our method against the DSTrust defense on the Fashion-MNIST dataset within a 50-client FL system, where 20% of clients are malicious. These tests were conducted under both IID and non-IID data environments. DSTrust effectively neutralizes the four baseline poisoning attacks when our method is not applied, highlighting its robustness as a defense mechanism. Despite DSTrust’s integration of both cosine similarity and distance metrics in its defense strategy, it fails to counteract the augmented attacks when our method is employed. Notably, our method achieves a maximum error rate increase of $0.521$, and an average error rate increase of $0.173$ across all $8$ test scenarios. These results demonstrate that merely understanding the adjustment strategies of our method, and subsequently integrating corresponding defense metrics, does not fundamentally negate the effectiveness of our method. Despite the adaptive defense’s attempt to incorporate both cosine similarity and distance metrics into DSTrust, it remains insufficient to thwart the enhanced capabilities of our method.
Distribution IID Non-IID
Attack FAvg FLT MKr Bulyan Median Trim DnC FLD Flame FAvg FLT MKr Bulyan Median Trim DnC FLD Flame
No Attack 0.319 0.328 0.338 0.324 0.330 0.337 0.315 0.334 0.336 0.336 0.358 0.330 0.336 0.338 0.336 0.347 0.326 0.374
Sign-Flipping 0.897 0.335 0.336 0.329 0.353 0.386 0.341 0.316 0.367 0.894 0.301 0.361 0.317 0.365 0.433 0.317 0.315 0.378
[…]
Trim Attack 0.431 0.323 0.422 0.428 0.434 0.432 0.340 0.339 0.362 0.423 0.345 0.469 0.429 0.474 0.451 0.352 0.299 0.355
[…]
Krum Attack 0.324 0.374 0.441 0.523 0.471 0.393 0.348 0.609 0.360 0.379 0.300 0.479 0.547 0.492 0.397 0.330 0.704 0.352
[…]
Table 14: Error rates on CIFAR-10 under cross-silo setting with 20% malicious clients and VGG-11 Net in the 30-client FL system.
Data Distribution IID Non-IID
No Attack 0.093 0.097 0.093 0.094 0.111 0.092 0.093 0.097 0.102 0.101 0.105 0.112 0.100 0.099
Sign-Flipping 0.893 0.157 0.092 0.106 0.110 0.108 0.100 0.963 0.122 0.104 0.103 0.118 0.114 0.106
[…]
Trim Attack 0.274 0.126 0.108 0.105 0.191 0.219 0.097 0.357 0.135 0.134 0.308 0.225 0.270 0.101
[…]
Table 15: Error rates on Fashion-MNIST under cross-silo setting with 20% malicious clients in a 100-client FL system.
Data Distribution IID Non-IID
Sign-Flipping 0.943 0.114 0.108 0.126 0.136 0.116 0.118 0.917 0.126 0.117 0.132 0.152 0.124 0.127
+ Neurotoxin 0.710 0.147 0.105 0.103 0.106 0.105 0.110 0.739 0.146 0.121 0.122 0.124 0.124 0.119
[…]
+ Neurotoxin 0.135 0.109 0.113 0.106 0.126 0.119 0.108 0.168 0.187 0.144 0.197 0.138 0.144 0.121
[…]
Table 16: Comparison with Neurotoxin under cross-silo setting on IID Fashion-MNIST in the 50-client FL system.
Appendix P Limitations and Future Work